The roots of the equation 25X2+ X +1 =0 are α^2 and β^2. Find the equation with

integral coefficients whose roots are 1/α and 1/β


1
Expert's answer
2021-09-12T23:52:43-0400
"25x^2+x+1=0"

By Viet Theorem


"\alpha^2+\beta^2=-\dfrac{1}{25}"

"\alpha^2\beta^2=\dfrac{1}{25}"

Then


"\dfrac{1}{\alpha^2}+\dfrac{1}{\beta^2}=\dfrac{\alpha^2+\beta^2}{\alpha^2\beta^2}"

"=\dfrac{-\dfrac{1}{25}}{\dfrac{1}{25}}=-1"


"\alpha\beta=-\dfrac{1}{5}\ \text{or}\ \alpha\beta=\dfrac{1}{5}"

"\alpha\beta=-\dfrac{1}{5}"


"\dfrac{1}{\alpha}(\dfrac{1}{\beta})=-5"

"(\dfrac{1}{\alpha}+\dfrac{1}{\beta})^2=\dfrac{1}{\alpha^2}+\dfrac{1}{\beta^2}+\dfrac{2}{\alpha\beta}"

"=-1+2(-5)=-11"

"\dfrac{1}{\alpha}+\dfrac{1}{\beta}=-i\sqrt{11}\text{ or }\dfrac{1}{\alpha}+\dfrac{1}{\beta}=i\sqrt{11}"

The quadratic equation is


"y^2+i\sqrt{11}y-5=0"

Or


"y^2-i\sqrt{11}y-5=0"


"\alpha\beta=\dfrac{1}{5}"

"\dfrac{1}{\alpha}(\dfrac{1}{\beta})=5"

"(\dfrac{1}{\alpha}+\dfrac{1}{\beta})^2=\dfrac{1}{\alpha^2}+\dfrac{1}{\beta^2}+\dfrac{2}{\alpha\beta}"

"=-1+2(5)=9"

"\dfrac{1}{\alpha}+\dfrac{1}{\beta}=-3\text{ or }\dfrac{1}{\alpha}+\dfrac{1}{\beta}=3"


The quadratic equation is


"y^2+3y+5=0"

Or


"y^2-3y+5=0"

"y^2+i\sqrt{11}y-5=0"


"y^2-i\sqrt{11}y-5=0"


"y^2+3y+5=0"


"y^2-3y+5=0"